A baseball coach shouts loudly at an umpire standing 3.3 meters away.

If the sound power produced by the coach is 3.0 × 10−3 W, what is the intensity of the sound when it reaches the umpire?

Answer in units of W/m2.
Answer #1

Solution:

Intensity = P / 4pi r^2

=> I = 3 * 10^-3 /(4 * 3.14 * 3.3 * 3.3)

=> I = 21.9 * 10^-6

=> I = 2.2 * 10^-5 W/m^2

Thanks